Estimate the temperature at which a reaction will
Answer #1

At equilibrium, Gibbs free energy change ΔG = 0

We know, according to Gibbs-Helmholtz equation, ΔG = ΔH- TΔS = 0

Given, ΔH = -10kJ, ΔS = -25J/K = -0.025 kJ/K

ΔH- TΔS = 0

i.e. -10kJ - T * (-0.025 kJ/K) = 0

i.e. T *  -0.025 kJ/K = -10 kJ

Hence, T = 400 K i.e. at 400 K temperature the reaction will be at equilibrium.